2x-y=6
step by step to fill in the missing values in a table containing
I am getting lost in the subtraction portion and addition.
X 0 _ 2 -2 4 -4
Y _ 0 _ _ _ _

Answer by stanbon(48575)

2x-y=6
Solve for "y":
Add y to both sides to get:
2x = 6+y
Subtract 6 from both sides to get:
y = 2x-6
----------------------
step by step to fill in the missing values in a table containing
I am getting lost in the subtraction portion and addition.
X 0 _ 2 -2 4 -4
Y _ 0 _ _ _ _
=====================
If x = 0, y = 2*0-6 = -6
If y = 0, 2x-6 = 0 ; 2x = 6 ; x = 3
If x = 2, y = 2*2-6 = 4-6 = -2
If x = -2, y= 2*-2-6 = -4-6 = -10
If x = 4, y = 2*4-6 = 8-6 = 2
If x = -4, y = 2*-4-6 = -8-6 = -14
=======================================
Cheers,
Stan H.